OpenStudy (anonymous):

A square ABCD has the following coordinates: A (1, 4), B (4, 4), C (4, 1), and D (1, 1). It is reflected over the y-axis. What are the coordinates of A´, B´, C´ and D´?

OpenStudy (anonymous):

(-1,4), (-4,4), (-4,1) and (-1,1) respectively

OpenStudy (anonymous):

A=1,-4 B=4,-4 C=4,-1 D=1,-1

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Did I get it wrong?

OpenStudy (anonymous):

Oh yes I am wrong

OpenStudy (anonymous):

its ok rid

OpenStudy (anonymous):

The x values are reflected inthe y axis

OpenStudy (anonymous):

gr8 rid and moorsy i m fan of u both..........!!!!!!